OpenStudy (anonymous):

if the lim x->0 f(x)/(x^2) = 8 then lim x->0 f(x) = and lim x->0 f(x)/x =

OpenStudy (watchmath):

The limit of \(f(x)\) as \(x\to 0\) must be zero. Otherwise the limit of \(f(x)/x^2\) is undefined. \(\lim f(x)/x=\lim f(x)/x^2\lim x=8\cdot 0=0\) (all the limit approaches 0)

OpenStudy (anonymous):

or in other words lim x->0 f(x)=8*x^2=0 and lim x->0 f(x)/x= lim(8*x^2)/x =lim 8x =0, as x->0